L. \underline{Default Parameter Values}

The default parameter values can be found in the Fortran block data files blkdata.F and kmixbdta.F. Within the job shell script a file model.data is created which contains a subset of the model parameters and these values override the default values in the block data files. Here we list some of the most important model parameters and their default values for the 2×2 and 3×3 configurations. Many of these can be changed at run time by specifying their values in the namelists which are defined in the file nldefs.h.

Stand-alone 2×2 model configuration (accelerated run)

Horizontal viscosity  am = 8.0 e+8
Background vertical viscosity  fkpm = 10.0
Horizontal diffusivity  ah = 0.0
Background vertical diffusivity  fkph = 0.3
Bottom friction  cdbot = 0.001
Eddy-induced advection coefficient  athkdf = 0.6 e+7
Isopycnal diffusivity  ahisop = 0.6 e+7
Inverse maximum slope  slmxr = 100.0
KPP critical bulk Richardson number  Ricr = 0.3
KPP critical local Richardson number  Riinfty = 0.7
KPP shear vertical viscosity  vvcric = 50.0
KPP shear vertical diffusivity  vdcric = 50.0
KPP convective vertical viscosity  vvclim = 1000.0
KPP convective vertical diffusivity  vdclim = 1000.0
Strong heat restoring under ice  restQp = 386.0
Strong salt restoring under ice  restFp = 2.77
Weak ocean heat restoring  restQm = 0.0
Weak ocean salt restoring  restFm = 0.092
Number of layers which can form ice  kmxice = 1
Tracer timestep  dtts = 24000.0
Baroclinic velocity timestep  dtuv = 2400.0
Barotropic velocity timestep  dtsf = 2400.0
Tracer vertical acceleration  dtxcel = 25∗1.0, 4∗5.0, 2∗6.0,
2∗7.0, 2∗8.0, 2∗9.0, 8∗10.0
Maximum number of scans  mxscan = 1000
Barotropic convergence criterion  crit = 2.0 e+7
Fraction of implicit Coriolis force  acor = 0.0

L-1 Coupled 2×2 model configuration (synchronous run)

The only changes from the 2×2 stand-alone parameter values are

Tracer timestep  dtts = 2400.0
Tracer vertical acceleration  dtxcel = 45∗1.0

and the four surface forcing restoring parameter values are irrelevant.

Stand-alone 3×3 model configuration (accelerated run)

Horizontal viscosity  am = 3.0 e+9
Background vertical viscosity  fkpm = 16.7
Horizontal diffusivity  ah = 0.0
Background vertical diffusivity  fkph = 0.5
Bottom friction  cdbot = 0.001
Eddy-induced advection coefficient  athkdf = 0.8 e+7
Isopycnal diffusivity  ahisop = 0.8 e+7
Inverse maximum slope  slmxr = 100.0
KPP critical bulk Richardson number  Ricr = 0.3
KPP critical local Richardson number  Riinfty = 0.7
KPP shear vertical viscosity  vvcric = 50.0
KPP shear vertical diffusivity  vdcric = 50.0
KPP convective vertical viscosity  vvclim = 1000.0
KPP convective vertical diffusivity  vdclim = 1000.0
Strong heat restoring under ice  restQp = 386.0
Strong salt restoring under ice  restFp = 2.77
Weak ocean heat restoring  restQm = 0.0
Weak ocean salt restoring  restFm = 0.023
Number of layers which can form ice  kmxice = 1
Tracer timestep  dtts = 35040.0
Baroclinic velocity timestep  dtuv = 3504.0
Barotropic velocity timestep  dtsf = 3504.0
Tracer vertical acceleration  dtxcel = 14∗1.0, 5∗5.0, 6∗10.0
Maximum number of scans  mxscan = 1000
Barotropic convergence criterion  crit = 1.0 e+6
Fraction of implicit Coriolis force  acor = 0.0

L-2